A person invested some amount at the rate of 12% simple interest and the remaining at 10%. He received yearly an interest of Rs.130. Had he interchanged the amounts invested ,he would have received an interest of Rs.134. How much money did he invest at different rates?
Answer by dkppathak(439)

A person invested some amount at the rate of 12% simple interest and the remaining at 10%. He received yearly an interest of Rs.130. Had he interchanged the amounts invested ,he would have received an interest of Rs.134. How much money did he invest at different rates?
let he invested Rs x at 12%
and Rs y for 10% as per given conditions
12% x + 10 %y = 130 (1)X5
10% x + 12% y = 134 (2)X6 by multiplying equation 1 and 2
12 x + 10 y = 13000 (1)X5
10 x + 12 y = 13400 (2)X6 by multiplying equation 1 and 2 as shown
60 X + 50 y = 65000
60 X + 72 Y = 80400 by subtracting
-22 Y = -15400
y = 15400/22 = rs 700
by substituting value of Y in equation 1
12 X +10 Y =13000
12 X + 10x 700=13000
12 X = 13000-7000
12 X = 6000
X = 6000/12
X= RS 500
Answer
Rs 500 0n 12% and Rs 700 on 10%